# Labelling Data with Unknown References

### v3

@@ 6 Experiments
We present two experiments: an empirical test of the theory (Section 6.1), and an application to a realistic scenario (Section 6.2). Ablation experiments are in Appendix D. We used two corpora: the data it knows how to label (in-phenomenon, or IP), and one it doesn’t (out-of-phenomenon; OOP). We use this naming to highlight the fact that the entire measurement is altered by the change in data–and, that, semantically, we are modelling different phenomena (i.e., distinct $f\colon X\rightarrow Y$). We refer to the hypothetical setup where we know the labels of a dataset as knowable (r. unknowable). Knowable scenarios are only to illustrate how the algorithm’s labelling adheres to the ‘true’ predictions, and in practice they would be unavailable. We run the No-Data Algorithm with $r=3$. See Appendix F for prompts and specifics on the models and methodology used.
@@ 6.1 Experiment 1: Synthetic Data without Natural Language
In this experiment we aimed to understand the behaviour of the No-Data Algorithm with a corpus close to the theoretical setup. We created two disjoint synthetic datasets over binary strings. This had the advantage of ablating out any potential memorisation concerns, and allowed us to control the rubrics, as shown in Rubric 1. For both OOP and IP, $\sigma$ is the majority vote. The test sets for both IP and OOP included 498 entries and are balanced.
@@ 6.1.1 Setup: Evaluators and Verifiers
Evaluators:
We used a decision tree (DT), and the highest-performing LLM we tested (o3-mini; OpenAI 2025a). [^2] The full results, in Appendix D, further include another closed- and two open-source models. Both observed a balanced training split of the IP dataset (DT was trained in it; the LLM observed it with five exemplars). We trained the DT to ensure that the evaluator learnt both IP’s $\sigma$ and $C$, even though during generation it only had access to $C$, and not $\sigma$. As ‘training’, the LLM observed the rubric in the prompt, and was made aware of the aggregation function only for the labelling step. For generation it was asked to pick the best candidate from the (training) dataset instead of outputting a new point, as we observed marked performance decreases otherwise (Appendix D). During OOP, the LLM was given the rubric for IP and an OOP datapoint. We tuned $\phi$ based on the knowable scenarios, to $\phi=0.4,0.1$ for the DT and the LLM, respectively.
Verifier:
We used a rule-matching algorithm that has access to the relevant rubrics (e.g., when testing IP, it observes the rubric for IP), implemented as described in Section 4.1.
@@ 6.1.2 Results
The results are in Table 1. The DT and the LLM adjusted to the predictions: accuracies with the No-Data Algorithm were within 2-2 and 0% of the knowable case in both IP and OOP. The success rate differences between IP (100% for DT; 81% for the LLM) and OOP (r. 5% and 28%) provided better information as to which setup was the one the evaluator truly knew. Of note, which we will cover in Section 7, is the number of flips performed by the algorithm in the DT OOP dataset: 46%.
We also performed ablation studies on the relationship of lying with evaluators and generators; the need for flipping the labels; and the effectiveness of generation strategies in various LLMs. Full details for all studies are available in Appendix D. We found that not flipping the labels led to lower accuracies in the No-Data Algorithm, especially when the evaluator lied and within IP. When the evaluator lied about knowing $\sigma$ or $f$, the success rates were lower in the IP case, but remained steady in the OOP case. The final study showed that LLM generation under this problem led to poor results in comparison to picking a string from a dataset.
@@
Table 1: Results for the known (test score if the labels were known) and unknown (output of the No-Data Algorithm) cases. The known case is a reference that in practice does not exist. For the unknown case, if we used the base generator, the performance is exactly the same as in the known case. The flip $\phi$ is set near the error rate in the known case ($0.4$ for the DT; $0.1$ for the LLM), since it is arguably the worst value it can have. This is needed in this experiment because the generator acts as an oracle, convincing the verifier every time, and hence the No-Data algorithm reduces to simple evaluation (the known case). However, the success rate is extremely low in DT OOP (unknown): even though there is no way for us to know whether the evaluator’s score is trustworthy, the success rate indicates deception.
@@ 6.2.1 Data
We created a multi-domain dataset (1,015 entries total) with randomly-drawn samples from OpenOrca (Lian et al., 2023), MMLU (Hendrycks et al., 2021), OpenCode Ahmad et al. (2025), and WildChat (Zhao et al., 2024). It was then professionally translated and annotated into West Frisian by four native speakers. See Section 9 and Appendix F for details on the corpus creation. West Frisian is a language that, albeit spoken by half a million people, is within the class of ‘exceptionally limited resources’ with ‘virtually no labelled data to use’, as per the taxonomy from Joshi et al. (2020).
Every entry in the corpus is a pair of prompts and outputs (LLM responses from GPT-4o), and the task was to judge whether the output is correct based on a rubric. The rubric, written in natural language, for IP consisted of six commonly-evaluated criteria ranging from ‘the response must be in West Frisian’ to ‘the response must be correct and without syntax errors’. For OOP we wrote a separate, almost-orthogonal rubric (e.g. ‘the response must be in Dutch’, as the model often responded in Dutch instead), and included a compound entry (‘one of B or C must be correct, otherwise the response is zero’), bringing the total criteria to seven. Full rubrics are in Appendix E. The $\sigma$ for this corpus was label 1 if all criteria were 1, otherwise 0, and performed the evaluation with the IP labels. Remark that this is equivalent to having a lying evaluator, in line with the setup from Section 6.1.
The evaluator and verifier was GPT-4.1 (OpenAI, 2025b). We split the dataset into train (500) and test (515). Ahead of running the experiments, we calibrated the prompt strategies that worked best under the IP known case for both the final labels and the per-criterion breakdown. While this is unrealistic (in practice users would not have access to this dataset), as before it allows us to have a ground truth from which to compare the No-Data Algorithm’s results.
Unlike the previous experiment, here neither the verifier nor the evaluator knew the label or the criteria values, and instead both created their own label sets. We observed that presenting exemplars to the model led to poor performance in both the IP and OOP cases, so the prompts do not use them.
@@ 6.2.2 Results
The results are in Table 2. They adjusted to the predictions, with high success rate and accuracies in the IP setting, and the converse in the OOP case. There were some nuanced differences when compared to the synthetic experiments from Section 6.1. Namely, the difference in performance was higher when $\phi$ was low (up to -3.9% and -2.2% accuracy in IP and OOP, r.). Nonetheless, success rates remained steady regardless of value of $\phi$. Additionally, we performed an ablation study comparing a model not proficient in West Frisian (Qwen 2.5B VL 7B; Qwen Team 2025). We found that in the IP case, where the model was given the correct rubric (i.e., it did not lie, but was wrong), the success rate was still low, hovering at around 35 in IP; and 1.8 for OOP (where it did lie), for all values of $\phi$. The full results are in Appendix D.3.
@@
Table 2: Performance of the LLM on the natural-language version of the No-Data Algorithm. We report the best prompts in the known case, which in turn informed the prompts we used in the unknown case. When using the predictions for the criteria to determine the label, the accuracy of the models is 80.0% (IP), and 50% (OOP).
@@ 7.1 Theory
Theorem 1 states that when the labeller has high accuracy, a low $\phi$ will be able to reproduce the label set. It is then possible to measure the successes to establish a confidence bound on the aggregate labels provided by the evaluator. It, however, has a different interpretation: when the accuracy of the evaluator is zero, there is no way to determine the labelling, regardless of $\phi$–it is completely unknowable. We argue that this makes sense: otherwise, it would be possible to create information out of thin air, and without any prior knowledge of $f$. Hence the primary purpose of the No-Data Algorithm is to establish trustworthiness, not to label data.
We assumed that $Y$ is a binary set, and that $X$ is the set of $k$-bit binary strings. The first assumption can be easily addressed by ‘binarising’ a set $Y$ with arity >2>2 by class-based separation. The second must be handled carefully from a mathematical point of view. For example, to model natural-language problems it would be inaccurate to assume that the $k$-bit encoding for $x\in X$ will retain correctness by solely working with the binary encoding of the strings. Extra information (e.g., semantics captured in a latent space) must be added to $x$, before proceeding as usual. Remark that, as per our experiments, this is not a concern empirically. Further developments of the theory addressing this and other aspects is in Appendix C.
@@ 7.2 Experiments
In both experiments, the results adjusted to the theory’s predictions: both the LLM and DT evaluators consistently had accuracies close to the known cases in IP/OOP, but distinct success rates in IP and OOP. This indicated that the No-Data Algorithm is robust to evaluator strategies attempting deception.
The high number of flips in the $\phi=0.5$ DT OOP case (Section 6.1) and $\phi=0.3$ LLM OOP (Section 6.2) support our interpretation of Theorem 1. Since the model’s accuracy is near-random in the OOP dataset–as indicated by the known case–the verifier will almost always catch a lie. Hence, most labels got flipped based on $\phi$–indeed, the exact proportion marked by $\phi$–and no information was gained from the labelling. It is worth noting that the success rate remained steady regardless of the value of $\phi$, even when the flips and accuracies didn’t. This means that it is an interpretable quantity, while $\phi$ is tuneable to attain ‘good’ labelled values.
Finally, it is worth noting that the evaluation of LLMs with LLMs, as per Section 6.2, depends strongly on the LLMs themselves. This is not a circular argument per se. The judge (of judges) is tuned on the IP case, and asked to evaluate an unkown model in the OOP case, as done in Section 6.2 and Appendix D.3, even when it had near-random performance in this task. This is because they are solving different tasks, and thus the accuracies did adjust to the known cases. From the perspective of future research directions, it does raise the question on what are the results measuring. We come back to this point in the next section.
@@ 8 Conclusion
We introduced an algorithm to determine trustworthiness of an evaluator in the case where labelled data does not exist. It works by performing randomly-selected checks evaluating specific parts of the datapoint to be labelled. We showed that these checks were sufficient and necessary to determine (w.h.p.) whether the evaluator knew the labels. Experimentally, the results adjusted to the theoretical predictions, with the No-Data Algorithm being able to determine lying evaluators under various scenarios, both synthetic and realistic. Of note was when the evaluator was randomly guessing the labels: while its accuracy was deceptively close to that of its known case, the success rate of the No-Data Algorithm was low, thus flagging that said labelling could not be trusted. Since this quantity remained steady across all values of $\phi$, it makes it a reliable metric to determine trust without needing tuning.
@@
One implication of the No-Data Algorithm’s correctness proof is that, when the evaluator cannot be fully trusted, the data cannot be annotated. This scenario is unsolvable: there is no prior information on the data, and hence it would be effectively creating information out of thin air otherwise. This is why the algorithm’s focus is on the successes: if an evaluator has low accuracy, but high success rate, it is very likely it is not lying; just wrong (e.g., the IP case in the experiments). Alternatively, high accuracy but low success rate indicates deception (r. the OOP case). Hence, the primary purpose of this algorithm is to establish trustworthiness and is up to the user to decide (to its own margin of trust) whether to accept or reject the labelling.
It is well-known that LLMs are sensitive to prompting. Establishing trust in an LLM evaluator in our setup was equivalent to establishing trust in the prompt, not the model. While effective in the two experiments we ran, this nuance could require extensions to the No-Data Algorithm to determine (w.h.p.) whether it is the model or the prompt what is failing to convince the verifier. As a starting point, we note that our realistic experiment slightly departed from the theoretical work by not assuming that the criteria was known, and instead relied on the evaluator and verifier to generate their own judgement of both $C$ and $\tilde{y}$. Nonetheless, our work has shown that it is possible, theoretically and empirically, to judge an evaluator–such as an LLM-as-a-judge–up to formal (mathematical) guarantees of correctness. The other implication of our work is that the algorithm’s setup allows the verifier to ‘self-calibrate’, by generating its own datapoints, since it knows which challenges will appear. This has both theoretical and practical implications that we leave for future exploration.
9 Ethics
This work’s contributions involve the ability to establish trust in an LLM. We are unaware of any potential misuse of this algorithm, albeit we admit we are unable to cover all potential applications. We, however, argue that releasing the code publicly has more benefits than potential harms. For the data annotation, all annotators were contracted through an annotator services company, and compensated at a rate of 28.7 USD/hr. Annotators were encouraged to take breaks and emphasised quality over speed through the annotation instructions.
10 Reproducibility Statement
All code has been included as part of the supplementary materials. It will be released under the MIT licence. The data will be released on the three licences from the original works. Detailed methodology, included model versioning, is in Appendix F. Prompts are in Appendix F.1, and also in the supplementary materials. Whenever available, we set the temperature to zero to ensure further reproducibility. Given the tendency of closed-source models to update and phase out existing models, we also include experiments with open models in Appendix D. The theoretical work includes the assumptions in Sections 3 and 5, and detailed proofs in Appendix A.

### v4

@@ 6 Experiments
We present two experiments: an empirical test of the theory (Section 6.1), and an application to a realistic scenario (Section 6.2). Ablation experiments and extensions are in Appendices C.1 and D. We used two corpora: one which the evaluators know how to label (in-phenomenon, or IP), and one they do not (out-of-phenomenon; OOP). This naming is to highlight that the measurement is altered by changing the dataset–semantically we are modelling different phenomena (i.e., distinct $f\colon X\rightarrow Y$). We refer to the hypothetical setup where we know the labels as knowable (r. unknowable). Knowable setups are only to illustrate how the algorithm’s labelling adheres to the ‘true’ predictions. In practice, they would be unavailable. We run the No-Data Algorithm with $r=3$. Remark that algorithmic baselines for this scenario are challenging, since approaches such as ensembling do not provide formal bounds of correctness, nor account for lying evaluators. See Appendix D.5 for a baseline analysis, and Appendix E for prompts and methodology details.
@@ 6.1 Experiment 1: Synthetic Data without Natural Language
In this experiment we aim to understand the behaviour of the No-Data Algorithm in a setup close to that of the theory (binary strings). For this, we created two disjoint synthetic datasets over $\{0,1\}$. This allows us to ablate any potential memorisation concerns, and to control the rubrics. The rubrics are in Table 1. For both OOP and IP, $\sigma$ is the majority vote. The test sets for both IP and OOP included 498 entries and are balanced.
@@ 6.1.1 Setup: Evaluators and Verifiers
Evaluators: We used a decision tree (DT), and the highest-performing LLM we tested (o3-mini; OpenAI 2025a). [^4] The full results, in Appendix D, further include another closed- and two open-source models. Both observed a balanced training split of the IP dataset. That is, we trained the DT to ensure that the evaluator learnt both IP’s $\sigma$ and $C$, even though during generation it only had access to $C$, and not $\sigma$. The LLM, as ‘training’, observed the rubric in the prompt, plus five exemplars, and was made aware of the aggregation function only for the labelling step. For generation it was asked to pick the best candidate from the (training) dataset instead of outputting a new point, as we observed marked performance decreases otherwise (Appendix D). During OOP, the LLM was given the rubric for IP and an OOP datapoint. We tuned $\phi$ to $\phi=0.4,0.1$ for the DT and the LLM, respectively. These are the worst-case scenarios, since, as per Theorem 5.2, it is preferable to have $\phi$ close to the knowable accuracy.
Verifier: We used a rule-matching algorithm matching the test set to its appropriate rubric.
@@ 6.1.2 Results
The results are in Table 2. The DT and the LLM adjusted to the predictions: accuracies with the No-Data Algorithm were within 2-2 and 0% of the knowable case in both IP and OOP. The success rate differences between IP (100% for DT; 81% for the LLM) and OOP (r. 5% and 28%) provided better information as to which setup was the one the evaluator truly knew.
We also performed ablation studies on the relationship of lying with evaluators and generators; the need for flipping the labels; and the effectiveness of generation strategies in various LLMs. Full details for all studies are available in Appendix D. We found that not flipping the labels led to lower accuracies in the No-Data Algorithm, especially when the evaluator lied and within IP. When the evaluator lied about knowing $\sigma$ or $f$, the success rates were lower in the IP case, but remained steady in OOP. These results align with the expectations from Section 5. The final study showed that LLM generation in this problem was slightly lower-performing (-3%) when compared to picking a string from the dataset. This is expected for non-natural language problems, does not impact the results, and does not extend to natural language. We discuss this further in Section 7.
Table 2: Results for the known (test score if the labels were known) and unknown (output of the No-Data Algorithm) cases. The known case is a reference that in practice does not exist. The flip $\phi$ is set near the error rate in the known case ($0.4$ for the DT; $0.1$ for the LLM), since it is arguably the worst value it can have. This is needed in this experiment because the generator acts as an oracle, convincing the verifier every time, and hence the No-Data algorithm reduces to simple evaluation (the known case). However, the success rate is extremely low in DT OOP (unknown): even though there is no way for us to know whether the evaluator’s score is trustworthy, the success rate indicates deception.
@@ 6.2 Experiment 2: Low-Resource Language Labelling
This experiment evaluates a setup much closer to a realistic scenario. The goal is to evaluate a chatbot’s suitability for deployment in a low-resource language. Added to that, the rubric contains ambiguous criteria, and the criteria values are not known to the verifier or the evaluator. Both components are LLMs-as-judges.
@@ 6.2.1 Data
We created a multi-domain dataset (1,015 entries total) with randomly-drawn prompts from OpenOrca (Lian et al., 2023), MMLU (Hendrycks et al., 2021), OpenCode (Ahmad et al., 2025), and WildChat (Zhao et al., 2024). The task is to judge whether the output of an LLM (GPT-4o) to these prompts is correct based on a rubric. Hence, every entry in the corpus is a pair of prompts and outputs professionally translated into West Frisian and annotated by four native speakers. West Frisian is a language that, albeit spoken by half a million people, is within the class of ‘exceptionally limited resources’ with ‘virtually no labelled data to use’, as per the taxonomy from Joshi et al. (2020). See Section 9 and Appendix E for details on annotation.
The rubric is written in natural language and consists of six criteria standard for chatbots deployed in user-facing applications, ranging from ‘the response must be in West Frisian’ to ‘the response must be correct and without syntax errors’. For OOP we wrote a separate, almost-orthogonal rubric (e.g. ‘the response must be in Dutch’, as the model often responded in Dutch), and included a compound entry (‘one of B or C must be correct, otherwise the response is zero’), bringing the total OOP criteria to seven. Full rubrics are in Appendix E.1. The $\sigma$ for this corpus is 1 if all criteria are 1, otherwise 0.
The evaluator and verifier is GPT-4.1 (OpenAI, 2025b), although with different prompts, and thus different functions. We split the dataset into train (500) and test (515). Ahead of running the experiments, we calibrated the prompt strategies that worked best under the IP known case for both the final labels and the per-criterion breakdown. This is analogous to the setting where a user would tune their verifier prior to deployment on the unknown cases.
This experiment is by design more challenging than the one from Section 6.1. Here neither the verifier nor the evaluator know the label or the criteria values: instead, both create their own label sets, and generate their own datapoints. The rubric has also ambiguity present (e.g., the IP rubric requests that the model must always answer ‘even if its answer is wrong’; but also that the response ‘must be correct’), and only the OOP rubric is decomposable. The ambiguity was unintentional, although not uncommon for realistic scenarios. We discuss this further in Section 7.
@@ 6.2.2 Results
The results are in Table 3. The LLMs tested had high success rates and accuracies in the IP setting, and the converse in the OOP case. There were some nuanced differences when compared to the synthetic experiments from Section 6.1. Namely, the difference in performance was higher when $\phi$ was low (up to -3.9% and -2.2% accuracy in IP and OOP, r.). Nonetheless, as expected, success rates remained steady regardless of value of $\phi$.
To further probe the impact of LLM choice on the results we also tested an evaluator not proficient in West Frisian (Qwen 2.5B VL 7B; Qwen Team 2025). This would be an evaluator which does not lie, but will be wrong regardless. We found that in the IP case, when the model was given the correct rubric, the success rate was still low, hovering at around 35% in IP; and 2% for OOP (where it did lie), for all values of $\phi$. The full results are in Appendix D.4. We also probed the effects of ternary label classification setup on the same dataset (Appendix C.1). Our findings largely adjusted to our other results, with the No-Data Algorithm spotting lying evaluators effectively.
Table 3: Performance of the No-Data Algorithm with GPT-4.1 as evaluator and verifier, and the natural-language setup as dataset. It can be seen that the performance of the evaluator adjusts to the predictions, with near-random accuracy in OOP and close-to-known accuracy in IP. Remark that the success rate remains comparatively stable, regardless of the chosen value of $\phi$. Hence the most appropriate quantity to measure is the success rate, while $\phi$ may be tuned later for higher accuracy gains.
@@ 7.1 Theory
Theorem 5.2 states that when the labeller has high accuracy, a low $\phi$ will be able to reproduce the label set. It is then possible to measure the successes to establish a confidence bound on the aggregate labels provided by the evaluator. It, however, has a different interpretation: when the accuracy of the evaluator is zero, there is no way to determine the labelling, regardless of $\phi$–it is completely unknowable. We argue that this makes sense: otherwise, it would be possible to create information out of thin air, and without any prior knowledge of $f$. Hence the primary purpose of the No-Data Algorithm is to establish trustworthiness, not to label data.
Our requirement that $Y$ is binary is sufficient from a theoretical angle, but worth discussing further. Arities larger than $2$ may be addressed with class-based separation, at the expense of a linear-time increase in runtime. This is an upper bound, and we conjecture that there exist versions of the No-Data Algorithm with smaller overheads. We expand upon this and other extensions to the theory in Appendix C. There we also show that, experimentally, our algorithm retains its guarantees of correctness under this setup. We leave its theoretical characterisation for further work.
@@ 7.2 Experiments
In both experiments both the LLM and DT evaluators consistently had accuracies close to the known cases in IP/OOP, but distinct success rates in IP and OOP. This indicates that the No-Data Algorithm is robust to evaluator strategies attempting deception.
The large number of flips in the $\phi=0.5$ DT OOP case (Section 6.1) and $\phi=0.3$ LLM OOP (Section 6.2) support our interpretation of Theorem 5.2. Since the model’s accuracy is near-random in the OOP dataset, the verifier will almost always catch a lie. Hence, most labels were flipped and no information was gained from the labelling. It is worth noting that the success rate remained steady regardless of the value of $\phi$, even when the flips and accuracies didn’t. Thus, success rate is the main metric to measure: $\phi$ may be tuneable post-hoc to attain better performance.
In our synthetic data scenario (Section 6.1), we noted that LLMs struggled with generating data for that problem, and picking it from a set was more effective. However, this was not the case in the natural-language scenario (Section 6.2). We argue that this is because LLMs are primarily trained with natural language, and identification (classification) and generation are two distinct problems (Kleinberg & Mullainathan, 2024); and certain automata-based problems like ours are harder for LLMs (De Wynter, 2026). From the perspective of our work, however, how an evaluator generates a datapoint makes no difference, so long as it can prove that it knows what it is doing.
Finally, remark that the evaluation of LLMs with LLMs, as in Section 6.2, depends strongly on the LLMs themselves. This is not a circular argument: the verifier (i.e., the prompt) was tuned on the IP case, and asked to evaluate an unknown LLM (r. evaluation prompt) in the OOP case, as done in Section 6.2 and Appendix D.4, even when it had near-random performance in this task. This is because the evaluator and verifier are solving different tasks, and hence why the accuracies adjusted to the known cases. We provide experiments in Appendix D–and specifically, Appendix D.4–to confirm this with different LLMs. However, from the perspective of future research directions, it raises the question on what are the (literature’s) results measuring. We come back to this point in the next section.
7.3 The Need for a Rubric
A key limitation of the No-Data Algorithm is the requirement of a well-designed rubric, and its performance hinges on it. If the rubric or aggregator is poorly defined, confidence in the algorithm’s results may be misplaced, and its reliability bounded by rubric quality. In other words, the No-Data Algorithm measures what it is told to measure.
Having a rubric is a fundamental aspect of the scientific method: verifying a hypothesis $H_{0}$ requires measuring a set of experimental outcomes, which are then used to accept or reject $H_{0}$. For this, outcomes must be interpreted (i.e., evaluated as criteria) and then aggregated. Thus, the responsibility for designing an effective rubric depends on the scientist (user), not the algorithm. This requirement is not unique to our work, or even annotation in machine learning. For example, there is work on rubric-based learning in reinforcement learning (He et al., 2025; Shao et al., 2025; Huang et al., 2025; Masters et al., 2025); and in evaluation with (not of) LLMs-as-judges (Lei et al., 2025).
However, needing a rubric and characterising it are two distinct considerations. For the latter, there are two cases that merit special attention: decomposability, and ambiguity. For decomposability, note that the algorithm requires this property for the proofs, but not necessarily at runtime. For example, the experiments from Section 6.2 used a decomposable rubric only in the OOP case. Since the No-Data Algorithm is agnostic to the data labels ($\sigma$ is unknown), it worked as intended.
Regarding ambiguity, recall that the rubrics from Appendix E.1 were ambiguous, and the experiments nevertheless succeeded. This further illustrates our argument on $H_{0}$: the performance of the LLMs on the (human-annotated) data was measured with respect to the human labels, which themselves were based on the rubric. Both annotators and LLMs used the same rubric. Hence, the measurement performed in Section 6.2 is whether the evaluator could act as a labeller the same way a human would.
@@ 8 Conclusion
Trust in a measurement, although foundational to the scientific method, is difficult to determine when labelled data is scarce. In this work we introduced the No-Data Algorithm to formally establish trust in an evaluator when labelled data does not exist. This algorithm works by posing carefully-designed challenges, inspired by zero-knowledge proofs, to an evaluator. We proved mathematically and empirically that this method is sufficient to establish an evaluator’s trustworthiness, or, alternatively, to detect deception by either sycophantic or uninformed models.
Our experiments in synthetic and natural-language scenarios showed close agreement to the theory, as well as some robustness to realistic settings (e.g., ambiguous rubrics and undefined criteria). Ablation studies with other LLMs and label arities showed extensibility and reinforced the findings.
@@
In our theoretical work we noted that an implication of the No-Data Algorithm is that, when the evaluator cannot be fully trusted, the data cannot be annotated. This is because there is no prior information on the data, and hence it would be effectively creating information out of thin air. This is why the algorithm’s focus is on the success rate. If an evaluator has low accuracy, but high success rate, it is very likely uninformed (e.g., the IP case in the experiments). High accuracy but low success rate indicates deception (r. OOP). Hence, it is up to the user to decide (to their own margin of trust) whether to accept or reject the labelling with respect to the rubric they designed. As noted in Section 9, we urge caution when interpreting the algorithm’s outputs, as they depend on the rubric itself.
We argue that our work reframes how evaluator reliability can be measured in situations with scarcely-labelled data. By establishing formal measures of trust, it is now possible to perform rigorous evaluations of ‘hard’ domains such as low-resource languages and medicine. This is particularly important due to the debates on the reliability and deployment-readiness of LLMs-as-judges, especially in sensitive applications.
In terms of further work, there are two open questions remaining, not addressed here due to scope: (1) how can one fully evaluate a model’s (not a prompt’s) trustworthiness? This is particularly important for open-ended scenarios, such as dialogue. Here, the principles behind the EV protocol still apply, albeit the probabilistic guarantees and challenges would differ. Hence, (2) what practical adaptations are needed for other data types and output signatures? Mathematically the full problem is solvable by assuming a binary string, but empirically this is often not feasible.
Our work shows that it is possible to judge an evaluator’s–such as an LLM-as-a-judge’s– trustworthiness with formal guarantees of correctness. It also enables practitioners to transition from assumed evaluator competence, to provable trust, thus enabling the building of more robust and trustworthy AI systems.
9 Impact Statement
This work’s contributions involve the ability to establish trust in an LLM. A possible risk is over-reliance on this algorithm’s measures of trust, especially if the underlying rubric is ill-posed. Users should understand that the No-Data Algorithm is only as valid as the rubrics provided. We thus urge users to not correlate high success rates with LLM infallibility. For the data annotation, all annotators were contracted through an annotator services company, and compensated based on seniority, at a rate starting at 28.7 USD/hr. Annotators were encouraged to take breaks and emphasised quality over speed through the annotation instructions.